-2[7d] = -14
The bar is a straight line, and I don't know how to type it. It is called the
absolute bar.
THanks for your help!

Answer by gonzo(654)

just to the right of where you typed ] there should be another key that has a vertical line over a backslash.
the vertical line is the absolute value symbol |
the backslash looks like this: \
the | is upper case.
the \ is lower case.
it may not look like one single line. it might look like two smaller vertical lines separated by a space in the middle.
hit the key and it should be a solid line on your screen.
on my keyboard it's the last key on the right below the backspace key and above the return key.
-----
your problem states that -2 * |7d| = -14
divide both sides of the equation by -2.
equation becomes:
|7d| = 14
-----
since 7d is within the absolute value sign, then d can be either positive or negative. the absolute value of d will always be positive.
-----
id d is positive, then |7d| = 7d
if d is negative, then |7d| = -7d
-----
solve for when d is positive.
then solve for when d is negative.
-----
when d is positive,
7d = 7
d = 1
when d is negative,
-7d = 7
-d = 1
d = -1
-----
your answer is:
d +/-1
-----
to prove, substitute in original equation
first for when d is positive
d = 1
-2*|7*1| = -2*|7| = -2*(+7) = -14
next for when d is negative
d = -1
-2*|7*(-1)| = -2*|-7| = -2*(-(-7)) = -2*(+7) = -14
-----
whether or not d was +1, or -1, the answer was the same because the absolute value of a positive number is a positive number, and the absolute value of a negative number is a positive number.
-----
good luck with finding your | key.